### v1

# A. CamillettiAI weather from Euro-Atlantic regimes

Abstract
We present a non-linear AI-model designed to reconstruct monthly mean anomalies of the European temperature and precipitation based on the Euro-Atlantic Weather regimes (WR) indices. WR represent recurrent, quasi-stationary, and persistent states of the atmospheric circulation that exert considerable influence over the European weather, therefore offering an opportunity for sub-seasonal to seasonal forecasting. While much research has focused on studying the correlation and impacts of the WR on European weather, the estimation of ground-level climate variables, such as temperature and precipitation, from Euro-Atlantic WR remains largely unexplored and is currently limited to linear methods. The presented AI model can capture and introduce complex non-linearities in the relation between the WR indices, describing the state of the Euro-Atlantic atmospheric circulation and the corresponding surface temperature and precipitation anomalies in Europe. We discuss the AI-model performance in reconstructing the monthly mean two-meter temperature and total precipitation anomalies in the European winter and summer, also varying the number of WR used to describe the monthly atmospheric circulation. We assess the impact of errors on the WR indices in the reconstruction and show that a mean absolute relative error below 80% yields improved seasonal reconstruction compared to the ECMWF operational seasonal forecast system, SEAS5. As a demonstration of practical applicability, we evaluate the model using WR indices predicted by SEAS5, finding slightly better or comparable skill relative to the SEAS5 forecast itself. Our findings demonstrate that WR-based anomaly reconstruction, powered by AI tools, offers a promising pathway for sub-seasonal and seasonal forecasting.
keywords:
Euro-Atlantic regimes, seasonal prediction, artificial intelligence, circulation patterns, reconstruction
1 Introduction
Predicting atmospheric conditions for the upcoming season is crucial in various social and economic applications. Accurate seasonal forecasts can guide agricultural planning, allowing farmers to optimize crop selection, planting, and harvesting schedules, supporting food security and market stability [Ceglar and Toreti, 2021]. In the energy sector, such forecasts help anticipate demand fluctuations, aiding in efficient resource allocation and grid management [Dorrington et al., 2020, Buonocore et al., 2022, El-Azab et al., 2025]. Additionally, governments and disaster preparedness agencies can use reliable forecasts to mitigate the impacts of extreme weather events, such as droughts or floods, thereby reducing economic losses and protecting vulnerable communities [Alfieri et al., 2013, Samaniego et al., 2019]. Despite recent advances, forecasting European weather on long-term timescales remains challenging for both dynamical [Palmer et al., 2004, Mishra et al., 2019] and statistical methods [Krikken et al., 2024]. While the weather forecast of physics-based numerical weather prediction (NWP) is limited to about two weeks due to the intrinsic chaotic nature of the atmosphere [Lorenz, 1963], statistical methods can explicitly leverage slowly varying, persistent, and more predictable components of the Earth system. By Rossby wave propagation, these large-scale phenomena give rise to teleconnection patterns between distant regions of the Earth [Wallace and Gutzler, 1981, Ambrizzi et al., 1995], opening a window of opportunity to enlarge the forecast horizons. In the northern hemisphere, the main mode of atmospheric variability, represented by the North Atlantic Oscillation (NAO) [Hurrell et al., 2003], or the strongly related Artic Oscillation (AO), exerts influence over North America, Europe and part of Asia [Deser et al., 2017, Trigo et al., 2002], and is characterized by a typical persistence of 10 to 30 days [Feldstein, 2003, Rennert and Wallace, 2009, Roberts et al., 2023].
The existence of persistent modes of the atmospheric state suggests that large-scale atmospheric circulation can be systematically decomposed into a finite number of dominant states [Rossby, 1940]. Weather regimes (WR), first introduced by Rex [1951], are recurrent, quasi-stationary, and persistent states of atmospheric circulation [Hochman et al., 2021]. In the Euro-Atlantic sector, the WR are strongly related to the NAO [Michelangeli et al., 1995] and are known to exert considerable influence over the European weather [van der Wiel et al., 2019, Trigo and DaCamara, 2000, Plaut and Simonnet, 2001, Hall and Hanna, 2018, Simpson et al., 2024] [^1] Some of the works mentioned here refer to different large-scale circulation patterns than WR, but share similar ideas and comparable results.. Their persistent nature [see e.g. Strommen et al., 2023, Mukhin et al., 2024b, for a related discussion], often influenced by teleconnection patterns [Roberts et al., 2023, Michel and Rivière, 2011], opens up the possibility of extending the forecast horizon in Europe beyond the limit posed by the chaotic nature of the atmosphere [Ferranti et al., 2015, 2018]. This justifies the increasing interest in assessing the capability of NWP [Ferranti et al., 2018, Büeler et al., 2021] and statistical models [Nielsen et al., 2022, Hall et al., 2019, Chattopadhyay et al., 2020, Bommer et al., 2025] to forecast the transitions between large-scale circulation regimes in the sub-seasonal and seasonal timescales. However, while much research has focused on the study of the correlation and the impacts of the WR on the European weather, the estimation of ground-level climate variables, such as temperature and precipitation, from Euro-Atlantic WR remains largely unexplored and limited to linear methods, e.g. considering the mean surface weather variables associated with the dominant regime [Madonna et al., 2021, Roberts et al., 2023]. Additionally, Gerighausen et al. [2024] highlight the importance of intra-regime variability and suggest the use of continuous WR indices to capture the regimes ‘subflavours’ in place of the categorical classification commonly found in the literature [Madonna et al., 2021, Grams et al., 2017, Plaut and Simonnet, 2001, Trigo and DaCamara, 2000].
In this study, we present an Artificial Intelligence (AI) model designed to reconstruct anomalies in European temperature and precipitation based on knowledge of the WR indices. The model can capture and introduce complex non-linearities in the relationship between the WR indices, which describe the state of the Euro-Atlantic atmospheric circulation, and the corresponding ground temperature and precipitation anomalies in Europe. After presenting the datasets (Sec. 2) and defining the WR and the target fields (Sec. 3), we present the AI model in Sec. 4. We evaluate its accuracy in reconstructing seasonal mean temperature and total precipitation anomalies in winter and summer (Sec. 5.1), as well as the impact of changing the number of indices used to represent the atmospheric circulation (Sec. 5.2). Given the potential future applications in a real forecasting scenario, we discuss the effects of introducing errors in the WR indices, establishing a lower bound on the WR prediction accuracy required to outperform the European Centre for Medium-Range Weather Forecasts (ECMWF) operational seasonal forecast system, SEAS5 (Sec. 5.3). In other words, we utilize the presented AI model as a mapping from the WR indices to the anomalies of interest to address the following questions:
  • How does the model’s reconstruction performance degrade as the number of input indices is reduced?
  • To what level of accuracy must the seven WR indices be forecasted to outperform SEAS5 in predicting seasonal mean two-meter temperature and total precipitation?
We also test the accuracy of the AI model in reconstructing the targeted anomalies using the SEAS5 forecast of the winters and summers WR indices (Sec. 5.4) as a first evaluation of the model in a concrete forecasting scenario.
In summary, the AI model presented here offers a promising approach to AI-driven long-range weather forecasting by leveraging the predictability of large-scale atmospheric circulation and its relation with surface-level variables. The model demonstrates good reconstruction capability, as well as notable adaptability and stability in the presence of errors in the WR indices, underlining its potential for seasonal forecasting applications. While it is applied in this study to reconstruct two-meter temperature and total precipitation anomalies on a monthly timescale, the architecture is sufficiently flexible to be adapted to other regions and variables without modification. Additionally, the model’s relatively lightweight design enables efficient training, further supporting its practicality for broader implementation.
2 Datasets
2.1 ERA5 reanalysis
ERA5, the fifth generation of ECMWF global reanalysis [Hersbach et al., 2020], is the current state-of-the-art reanalysis and the most widely used data set in climate sciences. The ERA5 reanalysis spans years from 1940 onward, with a latency of 5 days and a Cartesian grid resolution of 0.25°, corresponding to approximately 31 km horizontal resolution at the equator. Using a bilinear interpolation, we upscaled the ERA5 data from the original 0.25°×0.25°Cartesian grid to a 1°×1°Cartesian grid.
The geopotential height (GPH) is one of the most widely used variables for studying teleconnections and decomposing atmospheric circulation into a set of dominant modes. We use the daily mean GPH at 500 hPa (Z500) in the Euro-Atlantic region (80°W to 40°E, 30°N to 90°N) to compute the WR and the related WR indices (see Sec. 3.1) used as input in our model. The work aims to reconstruct the monthly mean two-meter temperature and total precipitation anomalies in Europe (i.e., within the domain 20°W to 40°E, 30°N to 70°N). We use the daily mean two-meter temperature and daily total precipitation to compute the respective daily anomalies, which are then aggregated monthly (see Sec. 3.3).
2.2 Numerical seasonal forecast: SEAS5
SEAS5 is the fifth-generation seasonal forecasting system developed by ECMWF [Johnson et al., 2019]. It provides global, long-range forecasts up to seven months ahead, using an advanced coupled ocean-atmosphere model. The SEAS5 forecast and hindcast of monthly aggregated GPH, two-meter temperature, and precipitation with 25 ensemble members were retrieved from the Copernicus Data Store (CDS) in the months from January 1981 to December 2024 on a 1°×1°regular grid over the Euro-Atlantic region previously defined.
3 Pre-processing
The following section details the pre-processing steps employed to derive the input and target variables for the AI-model. We then describe the model architecture, training and evaluation in Sec. 4.
3.1 Model’s input: weather regimes indices
Figure 1: Cluster mean of the $Z500$ field (black isocontours) and corresponding average $Z500$ anomalies in winter (December, January, February), associated with the seven year-round Euro-Atlantic WR. The average is performed in the 1981-2010 climatology. The ”No Regime” refers to the mean atmospheric circulation when there is no WR attribution according to the index life-cycle criteria [Grams et al., 2017].
There are different methods of extracting low-frequency modes of variability of the atmospheric circulation [refer to Hannachi et al., 2017, Schlef et al., 2019, Mukhin et al., 2024a, Spuler et al., 2024, for some examples]. In this work, we use the method presented in Grams et al. [2017], a natural extension of the procedure discussed in Ferranti et al. [2015]. The following briefly recap the procedure for obtaining the seven-year-round WR [Grams et al., 2017, Büeler et al., 2021].
The daily $Z500$ field is post-processed at each latitude $\phi$ and longitude $\lambda$ of the Euro-Atlantic sector to obtain the standardized anomaly field, $\tilde{\Phi}(t,\lambda,\phi)$:
  1. 1.
    The climatology period is defined as the days from the first of January of 1981 to the end of 2010
  2. 2.
    The calendar day climatology is obtained by applying a 15-day moving window centered on each calendar day (±$\pm~{}7$ days) in the climatology period;
  3. 3.
    The calendar day climatology is subtracted from the $Z500$ field, obtaining the $Z500$ anomalies, $\Phi(t,\lambda,\phi)$;
  4. 4.
    The resulting field is 5 days low-pass filtered using a Savitzky–Golay filter;
  5. 5.
    A calendar day normalization is computed by applying to the $Z500$ calendar standard deviation a 15-day moving window centered on each calendar day (±$\pm~{}7$ days) in the climatology period, and then taking the spatial average on the full domain;
  6. 6.
    The $Z500$ anomaly field, $\Phi(t,\lambda,\phi)$, is divided by the calendar day normalization.
[…]
The WR index, $I_{\rm wr}$, is defined as the normalized projection of the daily standardized anomaly field, $\tilde{\Phi}(t,\lambda,\phi)$, onto each cluster mean field, $\tilde{\Phi}_{\rm wr}(\lambda,\phi)$, where wr = (AT, ZO, ST, AR, EuBL, ScBL, GL):
$I_{\rm wr}(t)=\frac{P_{\rm wr}(t)-\left\langle P_{\rm wr}(t)\right\rangle}{\sqrt{\left\langle\left(P_{\rm wr}(t)-\left\langle P_{\rm wr}(t)\right\rangle\right)^{2}\right\rangle}}~{},$ (1)
[…]
$P_{\rm wr}(t)=\frac{1}{\sum_{\lambda,\phi}\cos\phi}\sum_{\lambda,\phi}\tilde{\Phi}(t,\lambda,\phi)\tilde{\Phi}_{\rm wr}(\lambda,\phi)\cos\phi$ (2)
The method outlined here for deriving the seven WR differs slightly from that used in Grams et al. [2017] and Büeler et al. [2021]. Key differences are detailed in Appendix B.
We are interested in quantifying the performance of the AI model in reconstructing the targeted anomalies by changing the number of WR used as input. Therefore, in addition to the seven weather regimes, we use the same procedure described before for the winter (December, January, February - DJF) and summer (June, July, August - JJA) seasons separately, changing the number of clusters to four. The four weather regimes obtained are the ZO, the GL, the EuBL, and the AR in winter (Fig. A.1), and the negative phase of the NAO (NAO-), a blocking state similar to the ScBL, the ZO (often referred as Atlantic Low [Cassou et al., 2005]) and the AT, in summer (Fig. A.2). We also compute the NAO index [Hurrell et al., 2003] separately for the summer and winter seasons, as the principal component (PC) of the first EOF. The analysis is performed on the same Euro-Atlantic region and the same climatology period used to compute the WR.
Finally, we aggregate the $I_{\rm wr}$ monthly. Indeed, we use the monthly average of the seven or four $I_{\rm wr}$ as an input variable in our AI model (see Sec. 4).
3.2 Perturbed indices
As previously mentioned, it is valuable to assess the model’s performance when using perturbed indices that simulate forecasting errors. This reflects a realistic scenario in which the predicted indices are not perfectly accurate.
Let us indicate with $I_{\rm wr}^{\prime}$ the set of indices with errors. The mean absolute error (MAE) between $I_{\rm wr}$ and $I_{\rm wr}^{\prime}$ is a dimensionless number that depends on the normalization of the anomalies and the normalization of the WR index itself. We then define the mean absolute relative error (MARE) as $\langle|I_{\rm wr}-I_{\rm wr}^{\prime}|/|I_{\rm wr}|\rangle$, which quantifies the typical absolute relative error between the correct index $I_{\rm wr}$ and the perturbed index $I_{\rm wr}^{\prime}$. We want to produce a perturbed index $I_{\rm wr}^{\prime}$ for a given value of the MARE. For this purpose, let us define $I_{\rm wr}^{\prime}=(1+f_{\rm wr})I_{\rm wr}$, where $f_{\rm wr}\sim N(0,\sigma)$ are random variables sampled from a normal distribution with mean zero and standard deviation $\sigma$. It follows that the mean error is zero by construction. In other words, there is no bias in the $I_{\rm wr}^{\prime}$ error. The MARE is $\langle|f_{\rm wr}|\rangle=\sqrt{2/\pi}\sigma$. Fixing the distribution of $f_{\rm wr}$, that is, the value of $\sigma$, for a given value of the MARE is straightforward. Also note that, since $f_{\rm wr}$ and $I_{\rm wr}$ are independent, the MAE of the perturbed indices, $\langle|I_{\rm wr}-I_{\rm wr}^{\prime}|\rangle$, is simply given by $\rm{MARE}\times\langle|I_{\rm wr}|\rangle$. Unlike MARE, the MAE, therefore, depends directly on the statistical properties of the original indices.
For every value of the MARE from 0 to 160%, we compute 50 random realization of $I_{\rm wr}^{\prime}$ sampling from $I_{\rm wr}\times N\left(0,\sqrt{\pi/2}~{}{\rm{MARE}}\right)$. Each realization is then passed as an input to the AI model (see Sec. 4) to obtain an estimate of the monthly average two-meter temperature and total precipitation in Europe for the months from January 2011 to December 2024 .
3.3 Target fields: temperature and precipitation anomalies
The daily two-meter temperature and total precipitation anomalies are computed following the same procedure at points 1 to 6 described in Sec. 3.1. Following the same pre-processing, the relevant relations between WR and target anomalies are preserved, allowing the AI model to exploit them for the reconstruction. The daily variables are then aggregated monthly: the daily mean two-meter temperature anomaly is averaged every day in a month while the daily total precipitation anomaly is summed. The monthly average two-meter temperature and total precipitation anomalies are the target variables that the AI model is asked to reconstruct.
Note that we intentionally do not detrend the two-meter temperature data, allowing the AI model to potentially capture the anthropogenic trend.
3.4 SEAS5 bias-corrected forecast
[…]
The calendar day climatology and standard deviation of $Z500$, two-meter temperature and precipitation, derived from ERA5 data (see steps 1 to 5 in Sec. 3.1), are aggregated to monthly values to obtain the SEAS5 monthly anomalies for these variables. Explicitly, the monthly average of the calendar day climatology (steps 1 to 4) is subtracted from the SEAS5 bias-corrected monthly forecasts. The resulting anomalies are then divided by the monthly average of the standard deviation computed in step 5.
Despite the AI model is trained on the $I_{\rm wr}$ computed from ERA5 $Z500$ standardized anomalies, $\tilde{\Phi}(r,\lambda,\phi)$, it is interesting to evaluate its performance in reconstructing the monthly average two-meter temperature and total precipitation anomalies using the predicted $I_{\rm wr}$ computed from the bias-corrected SEAS5 $Z500$ forecast, $I_{\rm wr}^{\rm SEAS5}$. To obtain $I_{\rm wr}^{\rm SEAS5}$ we project the forecast of the bias-corrected monthly mean $Z500$ standardized anomalies for each ensemble member $n$ and lead-time month $l$, $\tilde{\Phi}^{\rm SEAS5}(t,\lambda,\phi,l,n)$, to the seven mean cluster centers computed from the ERA5 daily anomalies, $\tilde{\Phi}_{\rm wr}(\lambda,\phi)$, (see Sec. 3.1). The calculations follow closely Eqs. 1 and 2, with the additional dimensions $l$ and $n$.
4 AI-model
Figure 2: Schematic representation of the model. The input vector $x$ with the $n$ WR indices, the year, and the encoded month is passed as an input to the model. After each layer, the output $y$ is passed to a ReLU activation function $\max(0,y)$. The $N=28$ dimensional vector, output of the second linear ResNet block, is stacked pixel-by-pixel onto the static data (latitude and longitude grids, land-sea mask, and digital terrain model). The resulting $(N+4)\times n_{\lambda}\times n_{\phi}$ tensor, where $n_{\phi}$ and $n_{\lambda}$ are the number of latitude and longitude grid points, respectively, is the input of the first convolutional ResNet layer. The number of features is then reduced by two by each of the next convolutional ResNet blocks.
The AI model is designed to reconstruct the monthly average of the standardized two-meter temperature and total precipitation anomalies in Europe (Sec. 3.3) from the monthly average WR index, $I_{\rm wr}$ (Sec. 3.1).
4.1 Model architecture and training
Fig. 2 schematically describes the model architecture. The monthly average $I_{\rm wr}$, for each WR, is passed along with the corresponding encoded month and year as input to the model. The month is encoded in twelve categories before all the inputs are processed first by a linear layer and subsequently to a sequence of two ResNet blocks, which takes the 20-dimensional input (seven WR, the year, and the twelve-dimensional encoded month) and yields a vector of $N=28$ dimensions. The $N$-dimensional vector is stacked pixel-by-pixel with the two-dimensional static data: the latitude grid, the longitude grid, the land-sea mask, and the digital terrain model (DTM) of the European region with the same resolution as the output two-dimensional field. The resulting $(N+4)\times n_{\lambda}\times n_{\phi}$ tensor, where $n_{\phi}$ and $n_{\lambda}$ are the numbers of latitude and longitude grid points, respectively, is the input of the next convolutional ResNet blocks, which reduce the number of features to one, i.e., to a two-dimensional field representing the target anomalies.
The AI model is trained separately to reconstruct two-meter temperature and total precipitation anomalies. It learns to approximate the target anomalies by adjusting its internal $\approx~{}30$k parameters to minimize the mean squared error (MSE) between the predicted and observed anomaly fields. To properly assess model performance and its generalization ability, the dataset is divided into distinct training and test periods. The training set spans from January 1940 to December 2010 , including the climatological period used to compute anomalies and define the WR indices. The model’s performance can be slightly influenced by the order in which the input and output pairs of the test dataset are passed to the model. We use random ordering and repeat the training procedure six times to remove better the dependence of the model’s performance from the training data random ordering. The test set covers the months from January 2011 to December 2024 , and a portion of this period, from January 2011 to December 2018, is also used for validation. Training is halted when the validation MSE reaches a minimum to prevent overfitting. It is important to note that the evaluation period includes the validation dataset used during model training. While this could potentially introduce bias into the reconstruction skill estimates, this issue is addressed and quantitatively assessed in Appendix D, where we demonstrate that any resulting bias is negligible. We choose to present results over the extended test period to ensure statistical robustness and enhance the reliability of the evaluation through a larger sample size. The limited size of the training dataset (70 years) constrains the choice of model architecture, preventing the use of larger and more complex models. Such models, with their higher number of parameters, would not only be computationally inefficient but also more prone to overfitting in data-scarce settings.
4.2 Evaluation metrics
Figure 3: AI model spatial MAE (top), ACC (middle) and CE (bottom) in reconstructing the winters (first and third columns) and summers (second and fourth columns) mean two-meter temperature and total precipitation anomalies using seven WR indices in the test years (from January 2011 to 2024 . Hatched areas mark regions with correlations that are not significant at $p>0.05$. Black contours indicate regions with ACC (CE) higher than 0.5 (0.3).
After performing an average on the six reconstructed anomalies by each of the six trained models, the quality of the reconstructed fields is evaluated using the metrics described in the following paragraphs.
Mean Absolute Error (MAE)
is a commonly used metric to evaluate the accuracy of a model in predicting numerical values. It measures the average magnitude of errors by considering the absolute differences between actual and predicted values. Given a dataset with $n$ observations, where $y_{i}$ represents the actual values and $\hat{y}_{i}$ the predicted ones, MAE is computed as
${\rm{MAE}}=\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|~{}.$ (3)
A lower MAE indicates a more accurate model. Since it relies on absolute differences, it treats overestimation and underestimation equally.
Anomaly Correlation Coefficient (ACC)
is a statistical measure to assess the similarity between predicted and observed anomalies in a dataset. It is commonly applied in climate science to evaluate deterministic forecast accuracy. ACC quantifies how well the anomalies correlate between predictions and observations. Given a set of anomalies for $n$ observations, where $y_{i}$ represents the observed anomalies and $\hat{y}_{i}$ the predicted ones, ACC is computed as
${\rm{ACC}}=\frac{\sum_{i=1}^{n}(y_{i}-\bar{y})\left(\hat{y}_{i}-\bar{\hat{y}}\right)}{\sqrt{\sum_{i=1}^{n}\left(y_{i}-\bar{y}\right)^{2}}\sqrt{\sum_{i=1}^{n}\left(\hat{y}_{i}-\bar{\hat{y}}\right)^{2}}}~{},$ (4)
where $\bar{y}$ and $\bar{\hat{y}}$ denote the mean values of observed and predicted anomalies, respectively. The ACC ranges from $-1$ to $1$, with values close to $1$ indicating a strong positive correlation, meaning that the model effectively captures the variations in anomalies. Conversely, values near 00 suggest no correlation, while negative values indicate an inverse relationship between predicted and observed anomalies.
To assess whether an observed correlation is statistically significant, we compute the associated p-value. The p-value represents the probability of obtaining higher or equal correlation assuming that there is no true correlation (i.e., under the null hypothesis of zero correlation). In this context, a p-value below 0.05 is typically considered statistically significant, indicating that there is less than a 5% probability that the observed correlation occurred by chance. Thus, correlations with $p<0.05$ can be interpreted as significant at the 95% confidence level.
Coefficient of efficiency (CE)
is a statistical measure used to evaluate the predictive skill of a model by comparing its performance to that of a validation climatology [Bürger, 2007], thus indicating the shared variance between the observed and estimated data throughout the verification period. Given $n$ observations, where $y_{i}$ represents the actual values, $\hat{y}_{i}$ the predicted values, and $\bar{y}$ the mean of observed values, the CE is computed for each grid point as,
${\rm{CE}}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y})^{2}}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}~{}.$ (5)
A value of $\rm{CE}=1$ indicates a perfect prediction, while $\rm{CE}=0$ suggests that the model performs no better than the validation climatology. Negative values imply that the model performs worse than the reference. In this study we use the climatology of the 2011 – 2024 testing period as validation climatology.
Since CE is not bounded from below, it is possible to have negative outliers in its distribution. We then prefer to use medians instead of averages when aggregating the skills discussed in this paragraph to avoid the outliers from significantly imbalancing the results.
To evaluate the aforementioned deterministic metrics on SEAS5, if no otherwise specified, we always use the ensemble mean as the predicted value.
5 Results
Figure 4: 2011 – 2024 winters (first and third columns) and summers (second and fourth columns) MAE (top row), ACC (middle row) and CE (bottom row) of the two-meter temperature anomalies and precipitation anomalies, reconstructed by the AI model trained with seven WR (red), four WR (orange), NAO (light-blue) and zero (blue) indices.
In the following sections, we evaluate the performance of the AI model in reconstructing monthly mean two-meter temperature and total precipitation anomalies in winter (DJF) and summer (JJA) from January 2011 to December 2024 . Sec. 5.1 presents the spatial reconstruction of these anomalies by the AI model trained with the seven year-round WR indices. We further investigate the model’s skill with respect to: (a) the number of WR indices used for training and as input (Sec. 5.2); (b) the impact of errors on $I_{\rm wr}$ on the reconstruction (Sec. 5.3); and (c) the use of $I_{\rm wr}^{\rm SEAS5}$ as input (Sec. 5.4). The first two aspects address how the reconstruction quality degrades, identifying a threshold to achieve a better seasonal forecast than SEAS5. The third point examines the accuracy of the reconstructed anomalies when using WR indices predicted by SEAS5. It can be interpreted as a hybrid NWP-AI framework for forecasting European mean two-meter temperature and total precipitation in winter and summer.
Although the AI model reconstructs anomalies over the entire European region as defined in Sec. 3.3, in the following analysis we restrict the evaluation to land areas only.
5.1 Anomaly reconstruction with seven WR
Fig. 3 shows the spatial MAE, ACC and CE in reconstructing the European monthly mean two-meter temperature and total precipitation anomaly using the seven-year-round WR indices, $I_{\rm wr}$, in the winters and summers during the test years.
For the two-meter temperature anomaly, the MAE maps indicate low reconstruction errors across most of Europe, especially in central and southern regions, with values frequently below 0.5. Errors are slightly higher in northern and eastern areas, particularly during winter, which is related to the higher variability and amplitude of the two-meter temperature anomalies in those regions. The ACC maps show high correlations (generally above 0.5) between reconstructed and observed anomalies, particularly over western and central Europe in both winter and northern east Europe in summer, but also in some part of the Mediterranean region. This indicates the model effectively captures temporal variability in these areas. The CE metric partially mirrors this pattern, with values approaching 0.3 in some areas, signifying that the model performs appreciably better than a climatological reference.
In the case of precipitation, the MAE values are more spatially variable and generally higher than for temperature, which reflects the inherently more chaotic and less predictable nature of precipitation. Nonetheless, several regions—especially parts of western Europe—exhibit relatively low MAE, particularly in winter. The ACC values for precipitation are moderate (typically between 0.4 and 0.7), with some improvement in winter relative to summer, suggesting that the WR-based model can partially capture precipitation variability, albeit with reduced skill compared to temperature. CE values for precipitation exhibit a similar spatial pattern, with positive skill in central Europe during winter but a marked decline in some southern and eastern regions during summer.
Summer extratropical variability is less influenced by the large amplitude stationary waves and tropical forcing that characterize the boreal winter climate. As a result, the summer weather in the Northern Hemisphere is more influenced by regional-scale processes, such as mesoscale convective weather systems and land-atmosphere coupling [Ambrizzi et al., 1995, Koster et al., 2000, Schubert et al., 2011]. This explains the lower skill reached by the AI model in summer compared to winter in all the skills tested. In particular, summer’s CE show that the model using the Euro-Atlantic WR is unable to appropriately reconstruct the mean two-meter temperature and total precipitation anomalies in the Mediterranean and Black Sea regions. For these areas, the reconstruction could be improved by including the Mediterranean large-scale circulation [Mastrantonas et al., 2021] and adding other predictors such as soil moisture or rainfall deficit indices [Vautard et al., 2007].
5.2 Impact of the number of WR on the reconstruction
In Fig. 4, the reconstruction skills (MAE, ACC, and CE) of the AI-model for monthly mean two-meter temperature and total precipitation anomalies during winter (DJF) and summer (JJA) are compared across configurations using seven WR indices (red), four WR indices (orange), the NAO index (light blue), and no indices (blue) as input. Analogously, Tab. 1 report the medians of the same metrics and the uncertainties on the last significant digit (reported in parenthesis), computed as the semi-difference of the skill metrics’ medians across multiple model initializations with different random seeds. The model trained without indices (No index) consistently shows the lowest skill scores for both temperature and precipitation. Despite this, it retains limited skill for two-meter temperature, with median ACC around $0.30\pm 0.07$ in winter and $0.32\pm 0.03$ in summer. This likely results from the AI-model capturing the long-term climate trend through the year and month inputs. However, the spatial distribution of this correlation lacks statistical significance in large regions, as shown in Fig. C.1. The inclusion of the NAO index, leads to some improvements. Since the NAO alone explains approximately 25% (16 %) of the atmospheric circulation variance in winter (summer), it significantly enhances the ACC for the precipitation in winter. Interestingly, for temperature, the skills are comparable between the models using the NAO index and without indices, indicating that much of the reconstruction skill for temperature likely arises from seasonal and anthropogenic trends rather than synoptic variability. Training the model with four WR indices yields substantial gains over the NAO model across both seasons and variables. For temperature, the typical improvements in ACC are about 29% in winter and 18% in summer, while for precipitation, they reach 100% (DJF) and 220% (JJA). The CE also improves, particularly for winter precipitation. Although CE remains near zero or slightly negative in JJA, particularly for precipitation, the four-index model exhibits far more dynamical skill and robustness than models with fewer indices. Expanding to seven WR indices results in smaller yet consistent improvements. The MAE remains largely comparable to the four WR setup, while ACC increases by 39% (DJF) and 20% (JJA) for temperature and by 15% (DJF) and 19% (JJA) for precipitation. The CE improves by approximately 260% for temperature and 41% for precipitation in winter but shows no substantial gains in summer, where skill levels remain low.
The gains between four and seven indices are strongly dependent on the season. While we do not observe substantial improvements in summer, using seven WR gives considerable enhancement in winter. We conclude that using at least four WR indices can provide an effective balance between reconstruction performance and index predictability, offering robust and meaningful reconstruction skills for both temperature and precipitation in winter and summer.
two-meter temperature
MAE ACC CE
7 WR 4 WR NAO No idx 7 WR 4 WR NAO No idx 7 WR 4 WR NAO No idx
DJF 0.29(1) 0.35(2) 0.36(1) 0.36 0.61(4) 0.44(8) 0.34(4) 0.30(7) 0.29(4) 0.08(6) 0.02(1) -0.06(3)
JJA 0.46(2) 0.46(1) 0.49(2) 0.54(2) 0.48(4) 0.40(4) 0.34(2) 0.32(3) 0.05(4) 0.03(2) -0.06(5) -0.17(4)
total precipitation
MAE ACC CE
DJF 0.29(1) 0.30(1) 0.36(2) 0.38 0.55(1) 0.48(1) 0.24(13) 0.05(4) 0.24(2) 0.17(1) -0.01(3) -0.04
JJA 0.36 0.37(1) 0.40(1) 0.41 0.38(1) 0.32(3) 0.10(5) 0.00(1) 0.08(2) 0.05(1) -0.04(1) -0.03
Table 1: Median of MAE, ACC, and CE for reconstructed monthly mean two-meter temperature and total precipitation during winter (DJF) and summer (JJA), using seven WR, four 4 WR, NAO, or no indices as input to the AI model. Uncertainties on the last significant digit (in parentheses) denote the semi-difference (half the difference between maximum and minimum) of the skill metrics’ medians across multiple model initializations with different random seeds. Uncertainties are not reported when different initialization seeds give consistent results.
5.3 Reconstruction degradation with indices errors
The 50 perturbed indexes for each value of the perturbation MARE are used as input to the AI model. The MAE, the ACC, and the CE for the summer and winter seasons can be evaluated as usual for each of the 50 reconstructed variables. After the skills are computed for each realization, we perform an average over the random realizations to evaluate the skills for the given MARE.
Figure 5: Spatial median of the mean two-meter temperature (red) and total precipitation (blue) ACC (triangles) and CE (rhomboids) reconstructed by the AI model in the winter seasons (top panel) and summers (bottom panel) from 2011 to 2024 , as a function of index MARE. The values are averaged over 50 random realizations of $I_{\rm wr}^{\prime}$ sampled from $I_{\rm wr}\times N\left(0,\sqrt{\pi/2}~{}{\rm{MARE}}\right)$. Horizontal lines are the SEAS5 spatial median of the ACC (solid) and CE (dashed) for the same variables and seasons.
The spatial median of the mean two-meter temperature and total precipitation ACC and CE, averaged over the 50 random realizations of the perturbed indices, are shown in Fig. 5 for the 2011 – 2024 winters and summers. Increasing the MARE on the indices, the AI model skills naturally degrade up to a point where they can be worse than the SEAS5 ensemble-mean ACC and CE. Both in winter and summer, the median ACC of the reconstructed mean two-meter temperature and total precipitation are always better than the SEAS5 ensemble-mean forecast ACC, indicating a strong efficacy of the AI model in capture the correlation between the Euro-Atlantic WR and European two-meter temperature and precipitation also in the presence of Gaussian distributed errors on the input indices. As for the ACC, the median CE of the mean two-meter temperature in winter and the precipitation in summer is always higher than the respective SEAS5 ensemble-mean median CE. More interesting is the case of the median CE of the reconstructed total precipitation in winter and mean two-meter temperature in summer, which becomes worse than the relative SEAS5 skill at MARE $\sim 80\%$ and $120\%$, respectively. Taking the lowest value of MARE of 80 % as a reference, this poses a lower bound on the forecast accuracy of the seven $I_{\rm wr}$ required to have a seasonal forecast of the mean two-meter temperature and total precipitation with higher deterministic skills than the bias-corrected SEAS5 ensemble mean forecast used in this work. We can express the results obtained here in terms of MAE of the perturbed indexes, $\langle|I_{\rm wr}-I_{\rm wr}^{\prime}|\rangle=0.35$ for MARE of 80 %.
5.4 AI-model in a concrete forecasting scenario
To test the AI model in a real forecasting scenario, we now analyze the reconstruction of the monthly mean two-meter temperature and total precipitation when using the monthly mean seven WR indices forecasted by the bias-corrected SEAS5, $I_{\rm wr}^{\rm SEAS5}$ (see Sec. 2.2), as model inputs. For both the AI model and SEAS5 the skills are computed for the ensemble mean, where the ensemble of the AI model reconstruction is obtained using as input the seven $I_{\rm wr}^{\rm SEAS5}$ forecast of each ensemble member. In the following, the SEAS5 forecasts for the winters (summers) are initialized on the first of November (May) to predict December, January and February (June, July and August).
Figure 6: MAE (top row), ACC (middle row) and CE (bottom row) for the SEAS5 (red) and the AI model (blue) using $I_{\rm wr}^{\rm SEAS5}$ as input. The first and second columns show the winter (left half) and summer (right half) skills for the two-meter temperature and precipitation, respectively.
Fig. 6 presents a comparative assessment of the hybrid NWP-AI model and the SEAS5 forecasting system for boreal winter and summer, evaluated using the usual metrics. The NWP-AI model typically exhibits better or comparable skills to the SEAS5 forecast. In particular, it achieves lower MAE values, particularly for the winter mean two-meter temperature and summer total precipitation anomalies. This enhanced performance is likely attributable to the use of a MSE loss during training, which suppresses variability in the model output. In contrast, SEAS5 tends to produce forecasts with higher variability, often resulting in higher MAE. The reduced MAE of the NWP-AI model translates into higher values of the CE, especially in cases where the discrepancy in MAE relative to SEAS5 is substantial. Nevertheless, both SEAS5 and NWP-AI model yield CE values that are often close to or below zero, indicating limited skill relative to the climatology of the test dataset. Regarding the ACC, the NWP-AI model slightly outperforms SEAS5 for the two-meter temperature in both seasons, but the ACC are comparable for the total precipitation anomalies.
6 Conclusions
This study presents and evaluates a lightweight AI-based model for reconstructing seasonal anomalies of two-meter temperature and total precipitation in Europe using Euro-Atlantic WR indices. The model demonstrates robust skill in capturing both spatial and temporal variability of temperature, with somewhat reduced performance for precipitation. Reconstruction skills are higher during winter, reflecting the stronger influence of large-scale circulation patterns on surface climate during this season. We test the model’s performance by varying the number of WR indices used as input. Despite the best skills being naturally reached by the models using seven WR, four indices represent a reasonable trade-off between reconstruction accuracy and forecast feasibility. We further assess the model’s robustness to errors in the input WR indices and identify a lower bound on their forecast accuracy necessary to outperform a state-of-the-art dynamical forecasting system, SEAS5. Specifically, to exceed SEAS5’s skill, the indices of winter (DJF) and summer (JJA) must be predicted with a MARE below approximately 80%, corresponding to a MAE of about 0.35. Finally, in a realistic forecasting setup where the WR indices are derived from bias-corrected SEAS5 forecasts, the hybrid NWP-AI model achieves comparable or superior performance relative to SEAS5, particularly for winter temperature and summer precipitation. These results highlight the potential of AI-driven approaches, grounded in dynamical predictors like WR indices, to complement traditional numerical forecasting methods for improved long-range climate prediction.
Nonetheless, several limitations remain. Training AI models usually require a large amount of data. We use monthly values from 1940 to 2010 , for a total of 852 points. The small training dataset forces the use of a lightweight model with a limited number of trainable parameters. Additionally, training with the MSE loss promotes smooth reconstructions with reduced variability. Other losses could be tested to have more realistic reconstructions. Moreover, the model is tested during the winters and summers in the 2011 – 2024 years. A small test dataset can significantly influence the statistical stability of the results presented in this work.
Despite these limitations, the approach is flexible and adaptable. The model can be extended to other regions, time scales, and surface variables. In particular, the sub-seasonal time scale presents a promising opportunity, offering both a richer dataset and potential for operational forecasting improvements. Recent studies have proposed new approaches for computing the WR [Paepe and Cruz, 2025, Spuler et al., 2024, Springer et al., 2024, Mukhin et al., 2024a]. Incorporating these WR could potentially enhance the reconstruction performance of the proposed AI-model.
The next efforts will be directed towards a fully AI-driven method to forecast the WR indices, which can be used as input to the AI model presented here to obtain a forecast of the surface-level variables.
[…]
The authors thank Ilaria Tessari and Federico Fabiano for the useful discussions on the year-round WR calculations. We also thank Leonardo Aragao for the confrontation on the bias-correction of SEAS5. This work was developed with financial support from ICSC–Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, Spoke4 - Earth and Climate, funded by European Union – NextGenerationEU.
Data availability
Daily post-processed ERA5 data on surface [Hersbach et al., 2024b] and pressure levels [Hersbach et al., 2023], was generated using Copernicus Climate Change Service information 2020, obtained from the Copernicus Climate Change Service [Hersbach et al., 2024a, b]. Neither the European Commission nor ECMWF is responsible for any use that may be made of the Copernicus information or data it contains. The AI-model implemented in this work are written in PyTorch Lightning [Falcon et al., 2020] and released under the MIT License through a GitHub repository (in preparation) based on the Lightning-Hydra-Template [Yadan, 2019]. The repository also contains all the script necessary to reproduce the results discussed in the paper (figures and tables). The data used to train the models will be made available on Zenodo (in preparation).
Appendix
Appendix A Four winter and summer WR
We compute the four WR for winter and summer using the same methodology described in Sec. 3.1, only changing the number of centroids used to cluster the data. The mean $Z500$ field (black isocontours) and the corresponding average $Z500$ anomalies during winter (December, January, February) and summer (June, July, August), associated with the four Euro-Atlantic WR, are presented in Fig. A.1 and Fig. A.2, respectively. The four WR are the ZO, ScBL, AT and GL. The two WR resemble the two phases of the NAO, i.e. the ZO and the GL.
Figure A.1: Cluster mean of the $Z500$ field (black isocontours) and corresponding average $Z500$ anomalies in winter (December, January, February), associated with the four Euro-Atlantic WR. The average is performed in the 1981-2010 climatology. The ”No Regime” refers to the mean atmospheric circulation when there is no WR attribution according to the index life-cycle criteria [Grams et al., 2017].
Figure A.2: Cluster mean of the $Z500$ field (black isocontours) and corresponding average $Z500$ anomalies in summer (June, July, August), associated with the four Euro-Atlantic WR. The average is performed in the 1981-2010 climatology. The ”No Regime” refers to the mean atmospheric circulation when there is no WR attribution according to the index life-cycle criteria [Grams et al., 2017].
Appendix B Differences in the WR computations with respect to Grams et al. [2017]
We now highlight the major differences to the procedure described in Grams et al. [2017] and in Büeler et al. [2021]. The authors of the mentioned work used the ERA-Interim reanalysis [Dee et al., 2011] six-hourly 500 hPa GPH from 11 January 1971 to 31 December 2015; we instead use ERA5 daily 500 hPa GPH from 1st of January 1940 to 31 December 2024 . Moreover, the anomalies computation, the EOF analysis, the K-means clustering, and the definition of the cluster’s mean field are performed only in the climatology period. Grams et al. [2017] applied 90- and 30-day moving window averages to smooth the calendar day mean and standard deviation. In contrast, we use a 15-day centered moving average for each calendar day. We observed that using window sizes larger than approximately 30 days tends to preserve seasonal trends in the resulting anomalies. This explains the minor difference that can be found when comparing Fig. 1 with the corresponding Supplementary Figure 1 in Grams et al. [2017] and Figure 1 in Büeler et al. [2021].
Appendix C Anomaly reconstruction without indices
Figure C.1: Spatial ACC (top) and CE (bottom) of the AI model reconstruction of the winter mean two-meter temperature and total precipitation anomalies without using the information encoded in the WR indices. The skills are computed in the test years (from January 2011 to 2024 . Hatched areas mark regions with correlations that are not significant at $p>0.05$. Black contours indicate regions with ACC higher than 0.5.
As a baseline for evaluating the added value of the WR indices, we trained the same AI model described in Sec. 4 to reconstruct monthly mean two-meter temperature and total precipitation anomalies without including the WR indices as input. The obtained spatial ACC and CE for the winter mean two-meter temperature and total precipitation anomalies are shown in Fig. C.1. Despite the absence of synoptic-scale information, the model exhibits low but positive ACC in certain European regions, likely due to its ability to learn and exploit the underlying climate trend present in the training data. Nevertheless, the CE value is around zero or negative, indicating that the model is not capable of replicating the amplitude of the fluctuations in the test dataset.
Appendix D Reconstruction excluding the validation dataset
In the main text, we presented the reconstruction skills the AI-model for the summers and winters from 2011 to 2024 , including the validation dataset form January 2011 to December 2018. The inclusion of the validation dataset is necessary to ensure a statistically relevant evaluation, as the independent test set alone is too limited in duration to provide robust skill estimates. The validation dataset is used to monitor the training and to stop it before overfitting occur, in order to guarantee optimal generalization of the AI-model. Therefore, including the validation dataset onto the test period can introduce a bias in the results, possibly improving the model’s performance. To assess the magnitude of this effect, we re-evaluate the reconstruction performance using only the independent test period, from January 2019 to 2024 .
Figure D.2: AI model spatial MAE (top), ACC (middle) and CE (bottom) in reconstructing the winters (first and third columns) and summers (second and fourth columns) mean two-meter temperature and total precipitation anomalies using seven WR indices in the test years (from January 2019 to 2024 . Hatched areas mark regions with correlations that are not significant at $p>0.05$. Black contours indicate regions with ACC (CE) higher than 0.5 (0.3).
Fig. D.2 shows the spatial distribution of the MAE, ACC and CE for the monthly mean two-meter temperature and total precipitation anomaly in 2019 – 2024, reconstructed using the seven year-round WR indices. This figure is the analogue of Fig. 3 and exhibits similar spatial patterns. As described in Sec. 5.1, model skill is consistently higher in winter and in central/western Europe, with temperature reconstruction being better than precipitation. The spatial patterns of high (low) ACC and CE (MAE) are largely shared between the two figures, though some quantitative differences are evident. In particular, where higher (lower) ACC and CE values are found, the AI-model tested on the shorter dataset scores better (worse) than the same AI-model tested on the extended dataset, especially for the two-meter temperature. The shorter validation period also limit the statistical significance of the correlations, as demonstrated by the large areas with p-value lower than 0.05 (less than 95% confidence). The p-value test is less reliable in discriminating significant from nonsignificant correlations in the regions with high or low ACC. or CE, which is computed as a temporal mean over the evaluation period (Eq. 5), using a shorter time window can exaggerate extreme values. This leads to larger differences in CE compared to the full-period evaluation, as reflected in Fig. 3.
two-meter temperature total precipitation
MAE ACC CE MAE ACC CE
winter 0.02 0.02 0.02 -0.01 -0.01 0.03
summer -0.04 0.04 0.16 -0.01 0.03 0.07
Table 2: Difference between medians of MAE, ACC and CE of the two-meter temperature and total precipitation reconstructions evaluated including the validation dataset (2011 - 2011 ) and excluding it (2019 - 2024 ). The skills are evaluated separately for winter (DJF) and summer (JJA). The model use the seven WR indices as input.
Tab. 2 presents the differences between the median skill metrics calculated with and without the inclusion of the validation dataset. The difference are typically small, of the order of few percent, indicating that including the validation data do not substantially bias in the results. The only notable increase occurs in the CE for summer temperature, likely because it is not bounded from below and degrades heavily when computed over fewer years. Excluding the validation data leads to more extreme negative CE values in certain regions, further amplifying discrepancies relative to the full dataset.
In conclusion, a comparison of model performance over the full evaluation period (2011–2024), which includes the validation years, and the independent test period (2019–2024) reveals reveals no consistent signs of overfitting or significant bias due to inclusion of the validation dataset.